OpenStudy (anonymous):

ln x = 1/3 (ln 16 + 2ln2)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ln16+2ln2 --> ln16 +ln2^2 -> ln16+ln4 -> ln16*4 \[ e^{lnx} = (1/3)ln(16*4)\] \[x=54e^{1/3}\]

OpenStudy (anonymous):

\[x=64e^{1/3}\]

OpenStudy (anonymous):

That's wrong, the answer is 4.

OpenStudy (anonymous):

Are you able to show me the work?

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

Google Chrome just crashed my openstudy Tab, oh lord.

OpenStudy (anonymous):

All that writing for naught.

OpenStudy (anonymous):

Essentially, you need two properties of logarithms: \[1.\log(a)+\log(b)=\log(a*b)\] \[2.b\log(a)=\log(a^b)\] See if you can work with that.

OpenStudy (anonymous):

Oh man :(. I just dont think I am writing it correctly and wanted to see how did you solve it

OpenStudy (anonymous):

so combine those two logarithms on the right to get \[log(16*4)\], which is \[log(4^3)\] and then use property "2" to get rid of that third power.

OpenStudy (anonymous):

What's property 2?

OpenStudy (anonymous):

The second thing I list in a previous post, scroll up.

OpenStudy (anonymous):

the answer is 4, correct

OpenStudy (anonymous):

I hate google chrome now.

OpenStudy (anonymous):

I'm sorry to bother you, but I'm just so confused. What about the 1/3?

OpenStudy (anonymous):

Well, now you have \[\frac{1}{3}\log(4^3)\] using that second property the coefficient in front of natural log becomes the power of 16, so \[\log(4^{3\frac{1}{3}})=\log(4)\]

OpenStudy (anonymous):

power of \[4^3\] not 16. Typo.

OpenStudy (anonymous):

thanks again! sorry about your computer problems :/